Compare the domain:

Domain of y= square root of x is 0 < x < infinity Domain of the inverse y=x2 is -infinity<x< infinity


What is the significant difference between the domains of y=square root of x and y= x2?


1
Expert's answer
2021-03-09T03:56:10-0500

Solution:

Domain of "y=\sqrt{x}" is "[0,\infty)"

Domain of "y=x^2" is "(-\infty,\infty)"

The significant difference between their domains is that "y=\sqrt{x}" cannot take negative values of "x" because the square root of negative value does not exist in the real world.

This can be seen in the following graph: